# Exploring the flow harmonic correlations via multi-particle symmetric and asymmetric cumulants in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV

### v3

@@ Abstract
We study multi-particle azimuthal correlations in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV. We use initial conditions obtained from a Monte-Carlo Glauber model and evolve them within a viscous relativistic hydrodynamics framework that eventually gives way to a transport model in the late hadronic stage of the evolution. We compute the multi-particle symmetric and asymmetric cumulants and present the results for their sensitivity to the shear and bulk viscosities during the hydrodynamic evolution. We also check their sensitivity to resonance decay and hadronic interactions. We demonstrate that while some of these observables are more sensitive to transport properties than traditional flow observables, others are less sensitive, making them suitable for studying different stages of the evolution.
@@ I Introduction
$\frac{dN}{d\phi}=\frac{1}{2\pi}\left[1+2\sum_{n=1}^{\infty}v_{n}cos\left[n\left(\phi-\psi_{n}\right)\right]\right]$ (1)
@@
which analytically relates multiparticle azimuthal correlators and flow degrees of freedom [31]. The average is calculated over all unique sets of $k$ different particles in a single event. This expression can be used to determine the properties of flow amplitudes $v_{n}$ and symmetry planes $\psi_{n}$ on an event-by-event basis. Apart from collective flow, other sources of correlations called non-flow are also present that typically involve only a subset of particles.
@@ I.1 Symmetric Cumulants of Flow Amplitudes
${\rm SC}(m,n)$$\equiv$ $\cos(m\varphi_{1}\!+\!n\varphi_{2}\!-\!m\varphi_{3}-\!n\varphi_{4})\right>\right>_{c}$ (4)
@@ I.2 Asymmetric Cumulants of Flow Amplitudes
These expressions are genuine multivariate cumulants. ${\rm AC}_{1,1}(m,n)$ corresponds to SC$(m,n)$, illustrating the generalisation aspect of the $ACs$. We can also normalize the $ACs$. This procedure has two benefits. Normalizing the results allows proper comparisons and determination of the initial state effects and the changes brought by the hydrodynamic evolution since the predictions for $ACs$ in the initial and final state do not have the same scale. Flow amplitudes depend on the transverse momentum, $p_{T}$, which leads to a similar dependence in any linear combinations of them, such as the $SCs$ and the $ACs$. The normalization eliminates this dependence and enables comparisons between models and data with different $p_{T}$ ranges [42]. Again, the normalization of the $ACs$ is done following the standard method from Ref. [37],
@@
In this paper, we have presented the measurement of ${\rm SC}(m,n)$ and ${\rm AC}_{2,1}(m,n)$ in Au+Au collisions at $\sqrt{s_{NN}}=200$ GeV using a hybrid hydrodynamic model. ${\rm SC}(m,n)$ have been computed in previous studies from hydrodynamic and transport models [43, 44, 42, 37, 45, 46, 47, 48], while ${\rm AC}_{2,1}(m,n)$ have been computed in transport models [46, 47]. ALICE collaboration has performed a systemic study of ${\rm SC}(m,n)$ and ${\rm NSC}(m,n)$ [49]. ALICE collaboration has also measured ${\rm AC}_{2,1}(m,n)$ for Pb+Pb collisions at $\sqrt{s_{NN}}=5.02\ TeV$ [50]. In addition to symmetric cumulants, various other correlators have also been studied [51, 52, 53, 54]. We show the sensitivity of these measurements to the transport properties, such as the shear viscosity to entropy density ratio, $\eta/s$, and the bulk viscosity to entropy density ratio, $\zeta/s$, of the medium.
@@ II Framework
$\eta/s$ $\zeta/s$
Parameter(Par.) Set I 0 0
Parameter(Par.) Set II 0.08 0
Parameter(Par.) Set III 0.08 $\zeta/s(T)$
@@
Particle list
Stage I Hydrodynamic evolution + particlization (Primordial)
Stage II Primordial + Resonace decays
Stage III Primordial + UrQMD
@@
In the following results, some observables will be presented by scaling with the average number of participants, $\langle N_{part}\rangle$. Table 4 lists the values of $\langle N_{part}\rangle$ in different centrality bins, which are determined using the Monte-Carlo Glauber model. The ${\rm SC}(m,n)$ and ${\rm AC}_{2,1}(m,n)$ observables have been computed using particles with transverse momentum ($p_{T}$) ranges of [0.2, 2.0] GeV and [0.01, 3.0] GeV, respectively.
@@
Table 4: $\langle N_{part}\rangle$ in the centrality classes from a Monte-Carlo Glauber model.
@@ III Error Estimation
The bootstrap method is used to efficiently estimate the error on ensemble-level observables using the Monte-Carlo approach, bypassing the complexities of standard error propagation. It involves randomly selecting events with replacement from an original sample to create bootstrap samples, allowing for the calculation of sampling variance.
Let $\hat{O}$ be the estimator of a statistic on which we intend to find the standard error. To estimate the standard error using the bootstrap method, follow these steps:
  • Given a parent sample of size $n$, construct $B$ independent bootstrap samples $X^{*}_{1},X^{*}_{2},\ldots,X^{*}_{B}$, each with $n$ data points randomly drawn with replacement.
  • Evaluate the estimator for each bootstrap sample.
@@ IV Results
We calibrated our model by using the produced particle yields, transverse momentum spectra and elliptic flow. We then compared these results to experimental data from the PHENIX and STAR Collaborations. The normalization factor for the system’s energy density was determined by matching the model and STAR data [70] charged hadron multiplicity $dN_{Ch}/d\eta$ in the 0-5% centrality bin. We also compared the $p_{T}$ spectra of identified particles in 0-5% centrality Au+Au collisions to STAR results [70]. Additionally, we compared the single-particle anisotropic flow coefficient $v_{2}$ of charged particles in 20-30% centrality Au+Au collisions with the PHENIX measurement [71]. The hydrodynamic evolution of the MCGM initial conditions, coupled to UrQMD, can reproduce the elliptic flow up to the mid-central collisions.
@@ IV.1.1 Four-particle symmetric cumulants
In this section, we present the results for ${\rm SC}(m,n)$ and ${\rm NSC}(m,n)$ obtained from our simulations. To evaluate their sensitivity to the hydro transport coefficients, we computed these observables using the three sets of ensembles listed in Table  2.
We found that both ${\rm SC}(2,3)$ and ${\rm SC}(2,4)$ are suppressed by the effects of shear and bulk viscosities. In Fig.[1], we illustrate how shear viscosity influences $v_{2}\{4\}$ and ${\rm SC}(2,3)$. For $v_{2}\{4\}$, the maximum change observed is about 10% in mid-central collisions. In contrast, ${\rm SC}(2,3)$ shows a change of approximately 40-50% across the centrality range. We also noted a similar dependence on the bulk viscosity of the medium, indicating that these observables are particularly sensitive to the transport coefficients.
The transport coefficients have more significant effects on ${\rm SC}(2,4)$ than on ${\rm SC}(2,3)$. A comparison of symmetric cumulants with $v_{3}\{4\}$ reveals a similar difference in their sensitivity to shear and bulk viscosities. It should be noted that $v_{n}\{4\}$ is more sensitive to the transport coefficients as $n$ increases [72]. This large sensitivity of symmetric cumulants primarily arises from the characteristics of anisotropic flow, as the symmetric cumulants involve higher powers of flow coefficients.
As discussed in the introduction, normalization removes the effect of the magnitudes of $v_{m}$ and $v_{n}$ on the value of the symmetric cumulant. Therefore, it should also eliminate the incidental sensitivity of ${\rm SC}(m,n)$ to shear and bulk viscosity. So, we also examined the sensitivities of the normalized cumulants. From Fig.[1], we observe that ${\rm NSC}(2,3)$ shows little sensitivity to shear viscosity. However, ${\rm NSC}(2,4)$ is more sensitive to both shear and bulk viscosities than $v_{2}\{4\}$.
@@ IV.1.2 Six-particle asymmetric cumulants
Now, we present the results for ${\rm AC}_{2,1}(m,n)$. In Fig. 1, we illustrate the effect of shear viscosity on $v_{2}\{4\}$ and ${\rm AC}_{2,1}(m,n)$. For ${\rm AC}_{2,1}(2,4)$, we observe a change of approximately 50% across the centrality range, although it is important to note that the 0-10% and 10-20% centrality classes contain large statistical uncertainties. We also find that the medium’s bulk viscosity similarly influences these observables.
The effects of the transport coefficients is found to be greater on ${\rm AC}_{2,1}(2,4)$ than on ${\rm AC}_{2,1}(2,3)$. Compared to traditional flow observables, these observables are more sensitive to transport coefficients. This magnified of asymmetric cumulants sensitivity primarily stems from the characteristics of anisotropic flow, as the asymmetric cumulants involve higher powers of flow.
As discussed in the introduction, normalization removes the effect of the magnitudes of $v_{m}$ and $v_{n}$ on the value of the asymmetric cumulant. Therefore, it should also eliminate the incidental sensitivity of ${\rm AC}_{2,1}(m,n)$ to shear and bulk viscosity. So, we also examined the sensitivities of the normalized asymmetric cumulants. From Fig.[1], it is evident that ${\rm NAC}_{2,1}(2,3)$ shows little sensitivity to shear viscosity. In contrast, ${\rm NAC}_{2,1}(2,4)$ demonstrates greater sensitivity to both shear and bulk viscosities compared to $v_{2}\{4\}$.
@@ IV.2.1 Four-particle symmetric cumulants
The hadronic afterburner plays a crucial role in interpreting data collected at RHIC [44]. To assess the sensitivity to hadronic interactions and resonance decay, we computed various observables at the end of the three stages of evolution mentioned in Table  3. We found that both ${\rm SC}(2,3)$ and ${\rm SC}(2,4)$ are suppressed by resonance decay and enhanced by hadronic interactions; however, the magnitudes of these effects differ between the two observables.
In Figs. [2][3][4], we illustrate how resonance decay and hadronic interactions impact $v_{2}\{4\}$ and ${\rm SC}(2,3)$. For $v_{2}\{4\}$, we observe a change of about 15% due to resonance decay, with an additional 5% change due to both hadronic interactions and resonance decay in mid-central collisions. In contrast, ${\rm SC}(2,3)$ shows a change of approximately 40% due to resonance decay across the entire centrality range and a change of 40-60% due to the combined effects of hadronic interactions and resonance decay. ${\rm SC}(2,4)$ is affected more strongly by resonance decay and late-stage hadronic interactions than ${\rm SC}(2,3)$. These observables, therefore, demonstrate higher sensitivity to resonance decay and hadronic interactions. The comparison of the symmetric cumulants with $v_{3}\{4\}$ also indicates similar differences in sensitivity to these effects. This increased sensitivity of symmetric cumulants arises mainly from anisotropic flow, with symmetric cumulants incorporating higher powers of flow.
As discussed in the introduction, normalization removes the effect of the magnitudes of $v_{m}$ and $v_{n}$ on the value of the symmetric cumulant. Therefore, it should also eliminate the incidental sensitivity of ${\rm SC}(m,n)$ to resonance decay and late-stage hadronic interactions. So, we also examined the sensitivities of the normalized symmetric cumulants. From Figs. [2][3][4], it is clear that ${\rm NSC}(2,3)$ shows little sensitivity to either resonance decay or late-stage hadronic interactions (UrQMD). In contrast, NSC(2, 4) exhibits greater sensitivity to both resonance decay and late-stage hadronic interactions compared to $v_{2}\{4\}$.
@@ IV.2.2 Six-particle asymmetric cumulants
Now, we present the results for ${\rm AC}_{2,1}(m,n)$. In Figs.[2][3][4], we show the effects of resonance decay and hadronic interactions on elliptic flow from the four-particle correlations and ${\rm AC}_{2,1}(2,3)$. For ${\rm AC}_{2,1}(2,3)$, we see a change of around 40% due to these in the mid-central collisions. The effects of resonance decay and late-stage hadronic interactions are found to be greater on ${\rm AC}_{2,1}(2,4)$ than on ${\rm AC}_{2,1}(2,3)$. Compared to the traditional flow observables, we see that these are more sensitive to resonance decay and hadronic interactions. This magnified sensitivity of asymmetric cumulants can be attributed to the properties of anisotropic flow, particularly as the asymmetric cumulants incorporate higher powers of flow.
As discussed in the introduction, normalization removes the effect of the magnitudes of $v_{m}$ and $v_{n}$ on the value of the symmetric cumulant. Therefore, it should also eliminate the incidental sensitivity of ${\rm SC}(m,n)$ to resonance decay and late-stage hadronic interactions. So, we also investigated the sensitivities of the normalized asymmetric cumulants. From Figures [2][3][4], it is evident that ${\rm NAC}_{2,1}(2,3)$ shows minimal sensitivity to both resonance decay and late-stage hadronic interactions (as represented by the UrQMD model). In contrast, ${\rm NAC}_{2,1}(2,4)$ demonstrates significantly greater sensitivity to these factors compared to $v_{2}\{4\}$. Therefore, these normalized cumulant measurements have the potential to differentiate between various transport models. Additionally, we examined the sensitivities of the normalized cumulants. From Figs. [2][3][4], it is clear that ${\rm NAC}_{2,1}(2,3)$ shows little sensitivity to either resonance decay or late-stage hadronic interactions (UrQMD). In contrast, ${\rm NAC}_{2,1}(2,4)$ exhibits greater sensitivity to both resonance decay and late-stage hadronic interactions compared to $v_{2}\{4\}$.
As a result, these normalized (a)symmetric cumulant measurements have the potential to differentiate between various models of quark-gluon plasma (QGP) evolution in hydrodynamic and transport models. ${\rm NSC}(2,3)$ and ${\rm NAC}_{2,1}(2,3)$, due to their insensitivity to hydro model parameters, resonance decay, and late-stage hadronic interactions, are particularly useful for constraining the initial state of the system’s evolution.

### v4

# Flow harmonic correlations via multi-particle symmetric and asymmetric cumulants in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV

@@ Abstract
We study multi-particle azimuthal correlations in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV. We use initial conditions obtained from a Monte Carlo Glauber model and evolve them within a viscous relativistic hydrodynamics framework that eventually gives way to a transport model in the late hadronic stage of the evolution. We compute the multi-particle symmetric and asymmetric cumulants and present the results for their sensitivity to the shear and bulk viscosities during the hydrodynamic evolution. We also check their sensitivity to resonance decay and hadronic interactions. We demonstrate that while some of these observables are more sensitive to transport properties than traditional flow observables, others are less sensitive, making them suitable for studying different stages of the evolution.
@@ I Introduction
$\frac{dN}{d\phi}=\frac{1}{2\pi}\left[1+2\sum_{n=1}^{\infty}v_{n}\cos\left[n\left(\phi-\psi_{n}\right)\right]\right]$ (1)
@@
which analytically relates multi-particle azimuthal correlators and flow degrees of freedom [31]. The average is calculated over all unique sets of $k$ different particles in a single event. This expression can be used to determine the properties of flow amplitudes $v_{n}$ and symmetry planes $\psi_{n}$ on an event-by-event basis. Apart from collective flow, other sources of correlations called non-flow are also present that typically involve only a subset of particles.
@@ I.1 Symmetric Cumulants of Flow Amplitudes
${\rm SC}(m,n)$$\equiv$ $\langle v_{m}^{2}v_{n}^{2}\rangle_{c}$
[…]
@@ I.2 Asymmetric Cumulants of Flow Amplitudes
These expressions are genuine multivariate cumulants. ${\rm AC}_{1,1}(m,n)$ corresponds to SC$(m,n)$, illustrating the generalization aspect of the $ACs$. We can also normalize the $ACs$. This procedure has two benefits. Normalizing the results allows proper comparisons and determination of the initial state effects and the changes brought by the hydrodynamic evolution since the predictions for $ACs$ in the initial and final state do not have the same scale. Flow amplitudes depend on the transverse momentum, $p_{T}$, which leads to a similar dependence in any linear combinations of them, such as the $SCs$ and the $ACs$. The normalization eliminates this dependence and enables comparisons between models and data with different $p_{T}$ ranges [41]. Again, the normalization of the $ACs$ is done following the standard method from Ref. [36],
@@
In this paper, we have computed ${\rm SC}(m,n)$ and ${\rm AC}_{2,1}(m,n)$ in Au+Au collisions at $\sqrt{s_{NN}}=200$ GeV using a hybrid model. ${\rm SC}(m,n)$ have been computed in previous studies from hydrodynamic and transport models [42, 43, 41, 36, 44, 45, 46, 47], while ${\rm AC}_{2,1}(m,n)$ have been computed in transport models [45, 46]. The ALICE collaboration has performed a systematic study of ${\rm SC}(m,n)$ and ${\rm NSC}(m,n)$ [48]. The ALICE collaboration has also measured ${\rm AC}_{2,1}(m,n)$ for Pb+Pb collisions at $\sqrt{s_{NN}}=5.02\ TeV$ [49]. In addition to symmetric cumulants, various other correlators have also been studied [50, 51, 52, 53]. We show the sensitivity of these measurements to the transport properties, such as the shear viscosity to entropy density ratio, $\eta/s$, and the bulk viscosity to entropy density ratio, $\zeta/s$, of the medium.
@@ II Framework
Parameter(Par.) Sets $\eta/s$ $\zeta/s$
I 0 0
II 0.08 0
III 0.08 $\zeta/s(T)$
@@
Stages Particle list
I Hydrodynamic evolution + particlization (Primordial)
II Primordial + Resonance decays
III Primordial + UrQMD
@@
In the following results, some observables will be presented by scaling with the average number of participants, $\langle N_{part}\rangle$. Table 4 lists the values of $\langle N_{part}\rangle$ in different centrality bins, which are determined using the Monte Carlo Glauber model. The ${\rm SC}(m,n)$ and ${\rm AC}_{2,1}(m,n)$ observables have been computed using particles with transverse momentum ($p_{T}$) ranges of [0.2, 2.0] GeV and [0.01, 3.0] GeV, respectively.
@@
Table 4: $\langle N_{part}\rangle$ in the centrality classes from a Monte Carlo Glauber model.
@@ III Error Estimation
The bootstrap method is used to efficiently estimate the error on ensemble-level observables using the Monte Carlo approach, bypassing the complexities of standard error propagation. It involves randomly selecting events with replacement from an original sample to create bootstrap samples, allowing for the calculation of sampling variance.
Let $\hat{O}$ be the estimator of a statistic on which we intend to find the standard error. To estimate the standard error using the bootstrap method, we follow these steps:
  • Given a parent sample of size $n$, we construct $B$ independent bootstrap samples $X^{*}_{1},X^{*}_{2},\ldots,X^{*}_{B}$, each with $n$ data points randomly drawn with replacement.
  • We evaluate the estimator for each bootstrap sample.
@@ IV Results
We calibrated our model by using the produced particle yields, transverse momentum spectra and elliptic flow. We then compared these results to experimental data from the PHENIX and the STAR Collaborations. The normalization factor for the system’s energy density was determined by matching the model and STAR data [69] charged hadron multiplicity $dN_{Ch}/d\eta$ in the 0-5% centrality bin. We also compared the $p_{T}$ spectra of identified particles in 0-5% centrality Au+Au collisions to STAR results [69]. Additionally, we compared the single-particle anisotropic flow coefficient $v_{2}$ of charged particles in 20-30% centrality Au+Au collisions with the PHENIX measurement [70]. The hydrodynamic evolution of the MCGM initial conditions, coupled to UrQMD, can reproduce the elliptic flow up to the mid-central collisions.
@@ IV.1.1 Four-Particle Symmetric Cumulants
In this section, we present our findings on the symmetric cumulants ${\rm SC}(m,n)$ and normalized symmetric cumulants ${\rm NSC}(m,n)$ obtained from our simulations. To evaluate their sensitivity to hydrodynamic transport coefficients, we analyzed these observables utilizing the three sets of ensembles detailed in Table 2.
Our results indicate that both ${\rm SC}(2,3)$ and ${\rm SC}(2,4)$ are significantly suppressed by the effects of shear and bulk viscosities. In Fig. 1, we illustrate the influence of shear viscosity on $v_{2}\{4\}$ and ${\rm SC}(2,3)$. The maximum observed variation for $v_{2}\{4\}$ is approximately 10% in mid-central collisions, while ${\rm SC}(2,3)$ displays a much larger change of around 40-50% across the entire centrality range. Although the comparison with $v_{3}\{4\}$ has not been shown, it yields similar results to the comparison with $v_{2}\{4\}$, with sensitivity increasing for higher $n$ values in $v_{n}$ [71]. This trend is similarly observed with respect to the bulk viscosity of the medium, underscoring the heightened sensitivity of these observables to the transport coefficients.
Furthermore, the impact of the transport coefficients on ${\rm SC}(2,4)$ is more pronounced than that on ${\rm SC}(2,3)$. This substantial sensitivity in symmetric cumulants arises primarily from the characteristics of anisotropic flow, as these cumulants incorporate higher powers of flow coefficients.
As discussed in the introduction, normalization serves to mitigate the influence of the magnitudes of $v_{m}$ and $v_{n}$ on the cumulants. Consequently, it is also expected to reduce the incidental sensitivity of ${\rm SC}(m,n)$ to shear and bulk viscosities. Therefore, we also examined the sensitivities of the normalized cumulants. From Fig. 1, it is apparent that ${\rm NSC}(2,3)$ exhibits minimal sensitivity to shear viscosity. However, ${\rm NSC}(2,4)$ shows greater sensitivity to both shear and bulk viscosities compared to $v_{2}\{4\}$.
@@ IV.1.2 Six-Particle Asymmetric Cumulants
Next, we present the results for the six-particle asymmetric cumulants ${\rm AC}_{2,1}(m,n)$. In Fig. 1, we also illustrate the effect of shear viscosity on ${\rm AC}_{2,1}(2,3)$. We found that for ${\rm AC}_{2,1}(2,3)$, the observed variation is around 40-60% across different centrality ranges; however, we note that the 0-10% and 10-20% centrality classes are subject to substantial statistical uncertainties. Additionally, the medium’s bulk viscosity similarly impacts these observables.
The influence of transport coefficients is notably more prominent for ${\rm AC}_{2,1}(2,4)$ compared to ${\rm AC}_{2,1}(2,3)$. Compared to traditional flow observables, asymmetrical cumulants display a heightened sensitivity to transport coefficients.
To further investigate these dynamics, we evaluated the sensitivities of the normalized asymmetric cumulants. From Fig. 1, it becomes evident that ${\rm NAC}_{2,1}(2,3)$ exhibits little sensitivity to shear viscosity. In contrast, ${\rm NAC}_{2,1}(2,4)$ displays a significantly greater sensitivity to both shear and bulk viscosities compared to $v_{2}\{4\}$.
@@ IV.2.1 Four-Particle Symmetric Cumulants
The hadronic afterburner is pivotal for interpreting data collected at the Relativistic Heavy Ion Collider (RHIC) [43]. To evaluate the sensitivity to late-stage hadronic interactions and resonance decay, we computed various observables at the conclusion of the three stages of evolution, as detailed in Table 3. Our analysis reveals that both ${\rm SC}(2,3)$ and ${\rm SC}(2,4)$ experience suppression due to resonance decay, while hadronic interactions lead to enhancement. Notably, the magnitudes of these effects exhibit significant differences between the two observables. Although the mean values show some consistency across different centralities, we exclude results for certain centralities for clarity, due to significant uncertainties.
Figs. [2], [3], and [4] illustrate the impact of resonance decay and hadronic interactions on $v_{2}\{4\}$ and ${\rm SC}(2,3)$. For $v_{2}\{4\}$, we observe a variation of approximately 15% due to resonance decay, and a 5% change attributable to the interplay between hadronic interactions and resonance decay in mid-central collisions. In contrast, ${\rm SC}(2,3)$ displays a change of roughly 40% driven by resonance decay across the full centrality range and a change of around 50% from the combined effects of hadronic interactions and resonance decay. Furthermore, ${\rm SC}(2,4)$ is found to be more significantly affected by hadronic interactions compared to ${\rm SC}(2,3)$. Consequently, these observables reveal enhanced sensitivity to resonance decay and hadronic interactions.
Next, we explore the sensitivities of the normalized symmetric cumulants, which are expected to reduce the incidental sensitivity of ${\rm SC}(m,n)$ to resonance decay and late-stage hadronic interactions. From Figs. [2], [3], and [4], it becomes apparent that ${\rm NSC}(2,3)$ shows minimal sensitivity to resonance decay or late-stage hadronic interactions (UrQMD). Conversely, ${\rm NSC}(2,4)$ demonstrates a heightened sensitivity to late-stage hadronic interactions compared to $v_{2}\{4\}$.
@@ IV.2.2 Six-Particle Asymmetric Cumulants
We now present the results for ${\rm AC}_{2,1}(m,n)$. Figs. [2], [3], and [4] display the influences of resonance decay and hadronic interactions on ${\rm AC}_{2,1}(m,n)$. For ${\rm AC}_{2,1}(2,3)$, we observe a noteworthy change due to these effects in mid-central collisions. The effects of late-stage hadronic interactions are observed to be more pronounced on ${\rm AC}_{2,1}(2,4)$ than on ${\rm AC}_{2,1}(2,3)$.
We also investigate the sensitivities of the normalized asymmetric cumulants. From Figs. [2], [3], and [4], it is evident that ${\rm NAC}_{2,1}(2,3)$ exhibits minimal sensitivity to either resonance decay or late-stage hadronic interactions (UrQMD). In contrast, ${\rm NAC}_{2,1}(2,4)$ illustrates substantially greater sensitivity to these factors compared to $v_{2}\{4\}$.
As a result, the normalized (a)symmetric cumulant measurements possess the capability to distinguish between diverse models of quark-gluon plasma (QGP) evolution, particularly within hydrodynamic and transport frameworks. ${\rm NSC}(2,3)$ and ${\rm NAC}_{2,1}(2,3)$, due to their insensitivity to hydrodynamic model parameters, resonance decay, and late-stage hadronic interactions, are especially valuable for constraining the initial conditions of the system’s evolution.